OpenStudy (anonymous):

Help with this math problem? Find 3 consecutive odd integers with a sum of 63

OpenStudy (anonymous):

Ok so in this problem we have 3 different integers that have to equal 63. Also they are 3 odd numbers. So for the first number we'll call it "x" . Now for the second number since it's odd we'll make it "x+2" Since x, will be odd we have to add it by another 2 to keep it odd and to make it consecutive. (odd #'s in a row) The third consecutive odd integer will be "x+4". Now we have: x+ (x+2)+(x+4)=63 [Add all like terms on left side of equal sign together] 3x+6=63= [Subtract 6 on both sides] 3x=57 [Divide both sides by 3] x=19. This means our first term of "x" is 19. Now for the 2nd term we have to do x+2, don't forget that "x" is 19, so we do : 19+2=21 for our 2nd term. Now for our last term we have to do "x+4" and 19+4=23. That means our three consecutive odd integers are 19,21,23. If u add them all up to double-check it u'll see it does in fact equal 63 :P So for a future problem u have 2: -Take a look at the word problem to see what it's saying (three consecutive odd integers equalling 63) -Place variables for the integers we don't know yet but r trying 2 find (x, x+2, x+4) -Find x (19) -Plug it back in for the actual terms :P x=19, x+2=19+2=21, & x+4=10+4=23) Did that help a bit 4 u 2 understand? :)

OpenStudy (anonymous):

Yes, thank you so much :D

OpenStudy (anonymous):

Np :P if u need any more help just ask and I"ll try 2 do my best :p sry it took so long tho lol stupid delete key on my keyboard stopped working -.- oh well :3

OpenStudy (anonymous):

Lol it's okay :D and thank you again.

OpenStudy (anonymous):

np :)